Can you, please, help me write this sequence of numbers in summation notation? Thank you! 12+4+4/3+4/9.
Answer by Boreal(15235)

sum (12/3^(n-1)), with n running from 1 to 4
n=1, value is 12/1
n=2, value is 12/3=4
n=3, value is 12/9=4/3
n=4, value is 12/27=4/9